# Relaxation and statistical equilibria in generalised two-dimensional flows

### v1

@@ 3.2 Spectral dynamics
At intermediate times, a power-law spectrum $E_{G}(k)\propto k^{-(\alpha+7)/3}$ emerges, which is indicated by the red dashed lines in Fig. 3. This scaling is consistent with the predictions for the dissipative generalised vorticity equations following K41 scaling arguments (Smithetal2002; jha2025cascades).
This Kolmogorov-like transient behaviour has also been identified in truncated 3D Euler and inviscid SQG (cichowlas2005effective; krstulovic2008two; krstulovic2009cascades; TeitelbaumMininni2012). During the relaxation to equilibrium a viscous-like inertial range develops through which thermalisation proceeds to lower wavenumbers. As time advances, thermalisation first develops at small scales and then progressively spreads towards larger scales, until the spectrum approaches the partial equilibrium state.
The analytical equilibrium spectra (14) derived in section 2 are plotted in Fig. 3 in dashed cyan color. As the large wavenumber modes thermalise, the numerical spectra start approaching the analytical spectra. Fig. 3 shows that at late times the large wavenumbers of the computed spectra are well predicted by the canonical ensemble theory derived in section 2, but the deviation is clearly noticeable in the small wavenumbers. This deviation, which is due to the partial thermalisation, has also been observed in previous studies (FoxOrszag1973; venaille2015violent) when $k_{c}/k_{\max}$ is small (Agouaetal2025). This is attributed to the fact that the canonical ensemble theory
@@ 3.3 Thermalisation
Here $k_{th}$ is a wavenumber, which is used to denote the modes above which thermalisation has taken place in the system, $k\geq k_{th}$, (TeitelbaumMininni2012). At long times we expect $E_{th}$ to reach a statistical steady state, whose average value will depend on $\alpha$. To remove this dependence, we subtract from the signal its long-time average, computed over the time interval $T$ defined by $0.9t_{\max}\leq t\leq t_{\max}$, where $t_{\max}$ denotes the maximum integration time. We denote the temporal average of $E_{th}$ as $\langle{E_{th}}\rangle_{T}$. The time $t_{\max}$ is chosen sufficiently large so that the average $\langle{E_{th}}\rangle_{T}$, taken over an interval spanning several tens of the turnover time $L/U$, is statistically converged.
@@
Now, we aim to understand the dependence of $\tau_{th}$ on $\alpha$, by plotting $\tau_{th}$ as a function of $\alpha$ in logarithmic scale in Fig. 4(b). To compare with the reduced model (27), we divide the $\tau_{th}$ obtained from the reduced model by $\alpha$, shown as dashed line in the plot. Figure 4(b) reveals a clear $\tau_{th}\propto 1/\alpha$ power-law scaling, indicating that $\tau_{th}$ diverges in the limit $\alpha\to 0$ and that the reduced model captures correctly this behaviour along with the expected prefactor. This validates the slow dynamics of the multiple scale analysis in Section 2.2, implying that the relevant time scale at which energy transfer occurs at small $\alpha$, scales like $1/\alpha$. The same scaling persists at a higher resolution of $512^{2}$ (see Fig. 4(b)) for $\alpha\leq 0.1$, where the wave number $k_{th}=30$ is kept the same, suggesting resolution-independent behaviour. The $1/\alpha$ scaling is expected to hold only for sufficiently small values of $\alpha$, and deviations naturally arise as $\alpha$ becomes order one. The partial thermalisation time increases with resolution for fixed $\alpha$ (see Fig. 4(b)), because a larger number of modes is required to equilibrate. This also makes the investigation of very small values of $\alpha$ computationally challenging, due to the rapid growth of $\tau_{th}$ at high resolution. Finally, we have also considered different initial conditions to test if they affect the $\tau_{th}\propto 1/\alpha$ scaling. In Appendix B, we show that the scaling does not depend on the initial conditions or on the specific choice of $k_{th}$, as long as a clear separation exists between the initially driven modes and the wavenumbers at which thermalisation is assessed.
The presence of a clear separation allows us to determine $\tau_{th}$ unambiguously by sharply identifying the onset of partial thermalisation in the system.
@@ 4 Conclusions
In this article, we focus on the canonical ensemble theory and the relaxation toward statistical equilibrium of generalised 2D flows. The parameter $\alpha$, which relates the generalised vorticity with the steamfunction controls the strength of the nonlinearity. We find that the statistical equilibrium solutions of the truncated system show $\alpha\mapsto-\alpha$ symmetry, where changing the sign of $\alpha$ is equivalent of exchanging $E_{G}\leftrightarrow\Omega_{G}$. In the dissipative system, changing the sign of $\alpha$ leads to a change in the direction of the cascade of the quantities $E_{G}$ and $\Omega_{G}$ (jha2025cascades). Hence, both the inviscid and the dissipative system show some aspects of the $\alpha\mapsto-\alpha$ symmetry.
@@
Although we understand reasonably well the scaling $\tau_{th}\propto 1/\alpha$, obtained for a specific set of initial conditions, we have not explored the full space of possible initial conditions. In this study, the simulations are initialised in regime I, where condensate formation occurs, leading the system into long-lived quasi-equilibrium states that remain far from the statistical equilibrium predicted from the canonical ensemble theory, as illustrated in Fig. 3 and reported in previous works (venaille2015violent; Agouaetal2025). This behaviour persists even as nonlinear interactions become progressively weaker in the limit of $\alpha\to 0$. By contrast, the initial conditions we tested in regimes II and III, but not reported here, relax toward statistical equilibrium states that are close to the canonical ensemble predictions. Both the reduced model and the numerical simulations are performed in the weakly nonlinear regime and at finite system size, under the condition $\alpha k_{\max}\ll 1$. A natural extension of the present study is to explore the thermodynamic limit $k_{\max}\gg 1$ as $\alpha\rightarrow 0$ with either $\alpha k_{\max}\to\text{constant}$ or $\alpha k_{\max}\to\infty$. Although the strength of individual nonlinear interactions vanishes in the limit $\alpha\to 0$, the number of interacting modes grows as $k_{\max}^{2}$, potentially compensating for the weakening of triadic interactions and giving rise to collective effects absent in finite-size systems. Such an interplay may lead to deviations from the $\tau_{th}\propto 1/\alpha$ scaling observed here and motivates the development of theoretical frameworks capable of addressing this regime. More broadly, while we have highlighted similarities between the dissipative and the inviscid system under the transformation $\alpha\rightarrow-\alpha$, it would be of considerable interest to investigate whether analogous symmetry principles arise in other systems that exhibit transitions in the direction of cascades of conserved quantities.

### v2

@@ 3.2 Spectral dynamics
At intermediate times, a power-law spectrum $E_{G}(k)\propto k^{-(\alpha+7)/3}$ emerges, which is indicated by the red dashed lines in Fig. 3. This scaling is consistent with the predictions for the dissipative generalised vorticity equations following Kolmogorov (1941) scaling arguments (Smith et al., 2002; Jha et al., 2025).
This Kolmogorov-like transient behaviour has also been identified in truncated 3D Euler and inviscid SQG (Cichowlas et al., 2005; Krstulovic and Brachet, 2008; Krstulovic et al., 2009; Teitelbaum and Mininni, 2012). During the relaxation to equilibrium a viscous-like inertial range develops through which thermalisation proceeds to lower wavenumbers. As time advances, thermalisation first develops at small scales and then progressively spreads towards larger scales, until the spectrum approaches the partial equilibrium state.
The analytical equilibrium spectra (14) derived in section 2 are plotted in Fig. 3 in dashed cyan color. As the large wavenumber modes thermalise, the numerical spectra start approaching the analytical spectra. Fig. 3 shows that at late times the large wavenumbers of the computed spectra are well predicted by the canonical ensemble theory derived in section 2, but the deviation is clearly noticeable in the small wavenumbers. This deviation, which is due to the partial thermalisation, has also been observed in previous studies (Fox and Orszag, 1973; Venaille et al., 2015) when $k_{c}/k_{\max}$ is small (Agoua et al., 2025). This is attributed to the fact that the canonical ensemble theory
@@ 3.3 Thermalisation
Here $k_{th}$ is a wavenumber, which is used to denote the modes above which thermalisation has taken place in the system, $k\geq k_{th}$, (Teitelbaum and Mininni, 2012). At long times we expect $E_{th}$ to reach a statistical steady state, whose average value will depend on $\alpha$. To remove this dependence, we subtract from the signal its long-time average, computed over the time interval $T$ defined by $0.9t_{\max}\leq t\leq t_{\max}$, where $t_{\max}$ denotes the maximum integration time. We denote the temporal average of $E_{th}$ as $\langle{E_{th}}\rangle_{T}$. The time $t_{\max}$ is chosen sufficiently large so that the average $\langle{E_{th}}\rangle_{T}$, taken over an interval spanning several tens of the turnover time $L/U$, is statistically converged.
@@
Now, we aim to understand the dependence of $\tau_{th}$ on $\alpha$, by plotting $\tau_{th}$ as a function of $\alpha$ in logarithmic scale in Fig. 4(b). To compare with the reduced model (27), we divide the $\tau_{th}$ obtained from the reduced model by $\alpha$, shown as dashed line in the plot. Figure 4(b) reveals a clear $\tau_{th}\propto 1/\alpha$ power-law scaling, indicating that $\tau_{th}$ diverges in the limit $\alpha\to 0$ and that the reduced model captures correctly this behaviour along with the expected prefactor. This validates the slow dynamics of the multiple scale analysis in Section 2.2, implying that the relevant time scale at which energy transfer occurs at small $\alpha$, scales like $1/\alpha$. The same scaling persists at a higher resolution of $512^{2}$ (see Fig. 4(b)) for $\alpha\leq 0.1$, where the wave number $k_{th}=30$ is kept the same, suggesting resolution-independent behaviour. The $1/\alpha$ scaling is expected to hold only for sufficiently small values of $\alpha$, and deviations naturally arise as $\alpha$ becomes order one. The partial thermalisation time increases with resolution for fixed $\alpha$ (see Fig. 4(b)), because a larger number of modes is required to equilibrate. This also makes the investigation of very small values of $\alpha$ computationally challenging, due to the rapid growth of $\tau_{th}$ at high resolution.
Finally, we have also considered different initial conditions to test if they affect the $\tau_{th}\propto 1/\alpha$ scaling. In Appendix B, we show that the scaling does not depend on the initial conditions or on the specific choice of $k_{th}$, as long as a clear separation exists between the initially driven modes and the wavenumbers at which thermalisation is assessed. The presence of a clear separation allows us to determine $\tau_{th}$ unambiguously by sharply identifying the onset of partial thermalisation in the system.
@@ 4 Conclusions
In this article, we focus on the canonical ensemble theory and the relaxation toward statistical equilibrium of generalised 2D flows. The parameter $\alpha$, which relates the generalised vorticity with the steamfunction controls the strength of the nonlinearity. We find that the statistical equilibrium solutions of the truncated system show $\alpha\mapsto-\alpha$ symmetry, where changing the sign of $\alpha$ is equivalent of exchanging $E_{G}\leftrightarrow\Omega_{G}$. In the dissipative system, changing the sign of $\alpha$ leads to a change in the direction of the cascade of the quantities $E_{G}$ and $\Omega_{G}$ (Jha et al., 2025). Hence, both the inviscid and the dissipative system show some aspects of the $\alpha\mapsto-\alpha$ symmetry.
@@
Although we understand reasonably well the scaling $\tau_{th}\propto 1/\alpha$, obtained for a specific set of initial conditions, we have not explored the full space of possible initial conditions. In this study, the simulations are initialised in regime I, where condensate formation occurs, leading the system into long-lived quasi-equilibrium states that remain far from the statistical equilibrium predicted from the canonical ensemble theory, as illustrated in Fig. 3 and reported in previous works (Venaille et al., 2015; Agoua et al., 2025). This behaviour persists even as nonlinear interactions become progressively weaker in the limit of $\alpha\to 0$. By contrast, the initial conditions we tested in regimes II and III, but not reported here, relax toward statistical equilibrium states that are close to the canonical ensemble predictions. Both the reduced model and the numerical simulations are performed in the weakly nonlinear regime and at finite system size, under the condition $\alpha k_{\max}\ll 1$. A natural extension of the present study is to explore the thermodynamic limit $k_{\max}\gg 1$ as $\alpha\rightarrow 0$ with either $\alpha k_{\max}\to\text{constant}$ or $\alpha k_{\max}\to\infty$. Although the strength of individual nonlinear interactions vanishes in the limit $\alpha\to 0$, the number of interacting modes grows as $k_{\max}^{2}$, potentially compensating for the weakening of triadic interactions and giving rise to collective effects absent in finite-size systems. Such an interplay may lead to deviations from the $\tau_{th}\propto 1/\alpha$ scaling observed here and motivates the development of theoretical frameworks capable of addressing this regime. More broadly, while we have highlighted similarities between the dissipative and the inviscid system under the transformation $\alpha\rightarrow-\alpha$, it would be of considerable interest to investigate whether analogous symmetry principles arise in other systems that exhibit transitions in the direction of cascades of conserved quantities.